# Near-Optimal Differentially Private $k$-Core Decomposition

### 12 Dec 2023 v1

@@ 3 Generalized Sparse Vector Technique
We introduce a multidimensional version of the AboveTreshold mechanism, which is the basis of the algorithms in the paper. In the standard AboveTreshold mechanism from Section 3.6 in [${}^{+}$14], we are given as input a threshold $T$ and an adaptive sequence of sensitivity $1$ queries $f_{1},f_{2},\ldots$, and the goal is to output the first query which exceeds the threshold. In the multidimensional version, we have a $d$-dimensional vector of thresholds $\vec{T}=(T_{1},\ldots,T_{d})$ and an adaptive sequence of sensitivity $\Delta$ vector-valued queries $\vec{f}_{1},\vec{f}_{2},\ldots$, and the goal is to output the first query for which the $j^{th}$ coordinate exceeds the threshold $T_{j}$, for each $j\in[d]$. We give a natural generalization of the 1-dimensional AboveThreshold mechanism in Algorithm 1 and give its privacy proof below.
Input: Private graph $G$, adaptive sequence of queries $\{\vec{f}_{i}\}$, threshold vector $\vec{T}$, privacy parameter $\varepsilon$, sensitivity $\Delta$.
@@
Observe that there are two sources of randomness in the algorithm: the noisy thresholds $\hat{T}_{j}$ for each $j\in[d]$ and the perturbations of the queries $\nu_{i,j}$ for each $j\in[d]$, $i\in[r(j)]$. We will fix the random variables $\nu_{i,j}$ for each $j\in[d]$ and $i<r(j)$ by simply conditioning on their randomness; for simplicity, we omit this from notation. The remaining random variables are $\hat{T}_{j}$ and $\nu_{r(j),j}$ for each $j\in[d]$; we will be taking probabilities over this randomness, but this will remain implicit in our notation. The main observation needed is that the output of $\mathcal{A}$ is uniquely determined by the iterations $r(j)$ of the first time we have $f_{i,j}(G)+\nu_{i,j}\geq\hat{T}_{j}$ for each $j\in[d]$. Thus, we can define the (deterministic) quantity
@@
$f_{r(j),j}(G)+\nu_{r(j),j}>\hat{T}_{j}\text{ and }\hat{T}_{j}\geq g_{j}(G)\text{ for each }j\in[d].$ (1)
@@
Since $G,G^{\prime}$ were arbitrary edge-neighboring graphs, this completes the proof.
@@ 4 Differentially Private $k$-Core Decomposition
In this section, we give our improved algorithm for differentially private $k$-core decomposition. We first present a variant of the classical (non-private) algorithm for $k$-core decomposition in Section 4.1, and then show how to make it private in Section 4.2. Finally, we show how to implement it in nearly-linear time in Section 4.3, while incurring only $(1+\eta)$-multiplicative error.
@@ 4.1 A Variation of the Classical Algorithm
Algorithm 2
@@ 4.2 Private Implementation of the Algorithm
11:           $\nu_{t}(v)\leftarrow\text{Lap}(8/\varepsilon)$
12:           if $d_{V_{t-1}}(v)+\nu_{t}(v)\leq k+\tilde{\ell}(v)$ then
@@
Algorithm 3
@@
For each iteration $t$, the $t^{th}$ query consists of the vector of $d_{V_{t-1}}(v)-k$ for each $v\in V$, where $V_{0}=V$ as in the algorithm; note that this matches with the queries on Line 12 of the algorithm. First, observe that the sensitivity of the vector containing $k-d_{V_{t-1}}(v)$ for each $v\in V$ is $\Delta=2$ for each iteration $t$, as needed in the algorithm. Next, observe that we can construct $V_{t}$ from $V_{t-1}$ using only the outputs $a_{t}(v)$ of the queries at the current iteration: given $v\in V_{t-1}$, we include $v$ in $V_{t}$ as well if and only if $a_{t}(v)=\bot$, which is equivalent to the condition in Line 12 of $d_{V_{t-1}}(v)+\nu_{t}(v)\leq k+\tilde{\ell}(v)$ [^2] Here, we implicitly assume that the randomness used by the algorithm and the AboveThreshold mechanism are the same. This is justified by conditioning on the random variables used by the two algorithms, proving differential privacy using the conditional probabilities. We then taking expectations over the conditional probabilities to conclude privacy for the full algorithm.. Thus, this is a feasible sequence of adaptive queries for the AboveThreshold mechanism.
@@
By the density function of the Laplace distribution, we know that we have $|\nu_{t}(u)|\leq\frac{40\log{n}}{\varepsilon}$ and |~$|\tilde{\ell}(u)|\leq\frac{20\log{n}}{\varepsilon}$ each with probability at least $1-\frac{1}{n^{5}}$. Since there are at most $O(n^{3})$ such random variables, taking a union bound over all nodes $u\in V$ and all iterations of the loops, we have the above guarantee with probability at least $1-{O}\left(\frac{1}{n^{2}}\right)$. We condition on the event that the above inequalities hold true for each $t\in[T]$ and each $u\in V$ for the remainder of the proof.
Fix an arbitrary iteration $k$ of the while loop starting on Line 7. Let $H$ be the set of nodes remaining in $V_{t}$ at the end of the while loop in Lines 8—16. We claim that ($i$) all nodes $u\in H$ have core number at least $k-\frac{60\log{n}}{\varepsilon}$ and ($ii$) all nodes $u\not\in H$ have core number at most $k+\frac{60\log{n}}{\varepsilon}$. To see ($i$), consider the subgraph $H$; we claim that each node $u\in H$ has induced degree at least $k-\frac{60\log{n}}{\varepsilon}$ in $H$. Indeed, since each node in $H$ was not removed in the final iteration of the while loop in Lines 8—16, we have that $d_{H_{t}}(u)+\nu_{t}(u)\geq k+\tilde{\ell}(u)$ for each $u\in H$. But since we have assumed $|\nu_{t}(u)|\leq\frac{40\log{n}}{\varepsilon}$ and |~$|\tilde{\ell}(u)|\leq\frac{20\log{n}}{\varepsilon}$, our desired bounds follow directly by the triangle inequality. To see ($ii$), let’s suppose for contradiction that the $k$-core value of $u$ is $k(u)>\ell+\frac{60\log{n}}{\varepsilon^{\prime}}$. Then there exists a subgraph $u\in K\subseteq V$ where the induced degree of each node $v\in K$ is $d_{K}(v)\geq k(u)$. But for such a subgraph $K$, the condition in Line 12 will always be true (again, because $|\nu_{t}(u)|\leq\frac{40\log{n}}{\varepsilon}$ and |~$|\tilde{\ell}(u)|\leq\frac{20\log{n}}{\varepsilon}$) so $K\subseteq H$. But since $u\in K$, this contradicts the fact that $u\not\in H$.
@@ 4.3 Efficient Implementation of the Algorithm
11:           $\nu_{t}(v)\leftarrow\text{Lap}(8/\varepsilon)$
12:           if $d_{V_{t-1}}(v)+\nu_{t}(v)\leq k+\tilde{\ell}(v)$ then
@@
13:     $t\leftarrow t+1$, $V_{t}\leftarrow V_{t-1}$//this is only notational; we only store the set $H_{t}$ for a single $t$
@@ 5.1 Differentially Private Densest Subgraph
Algorithm 6 outputs an $\varepsilon$-edge differentially private $\left(2,O\left(\frac{\log(n)}{\varepsilon}\right)\right)$-approximate densest subgraph with probability $1-O\left(\frac{1}{n^{c}}\right)$ for any constant $c\geq 1$.
@@ 5.2 Differentially Private Low Out-Degree Ordering
[${}^{+}$22] defines the low out-degree ordering problem to be finding a differentially ordering where the out-degree of any vertex is minimized when edges are oriented from vertices earlier in the ordering to later in the ordering. Specifically, there exists an ordering of the vertices of any input graph where the out-degree is at most the degeneracy $d$ of the graph, where the degeneracy of the graph is equal to the maximum $k$-core value. The previous best differentially private algorithm for the problem gives out-degree at most $(4+\eta)d+O\left(\frac{\log^{3}n}{\varepsilon}\right)$. In this paper, we give a novel algorithm that obtains out-degree bounded by $d+O\left(\frac{\log n}{\varepsilon}\right)$, with high probability.

### 29 Feb 2024 v2

@@ 3 Generalized Sparse Vector Technique
We introduce a simple, novel multidimensional generalization of the AboveThreshold mechanism, which is the basis of the algorithms in this paper. In the standard AboveThreshold mechanism from Section 3.6 in [${}^{+}$14], we are given as input a threshold $T$ and an adaptive sequence of sensitivity $1$ queries $f_{1},f_{2},\ldots$, and the goal is to output the first query which exceeds the threshold. In the multidimensional version, we have a $d$-dimensional vector of thresholds $\vec{T}=(T_{1},\ldots,T_{d})$ and an adaptive sequence of sensitivity $\Delta$ vector-valued queries $\vec{f}_{1},\vec{f}_{2},\ldots$, and the goal is to output the first query for which the $j^{th}$ coordinate exceeds the threshold $T_{j}$, for each $j\in[d]$. Here $\Delta$ is defined as the sensitivity for all vectors. We give a natural generalization of the 1-dimensional to the multidimensional AboveThreshold mechanism in Algorithm 1 and give its privacy proof below.
Input: Private graph $G$, adaptive queries $\{\vec{f}_{1},\dots,\vec{f}_{n}\}$, threshold vector $\vec{T}$, privacy $\varepsilon$, sensitivity $\Delta$.
Output: A sequence of responses $\{\vec{a}_{1},\dots,\vec{a}_{n}\}$ where $a_{i,j}$ indicates if $f_{i,j}(G)\geq\vec{T}_{j}$
@@
Observe that there are two sources of randomness in the algorithm: the noisy thresholds $\hat{T}_{j}$ for each $j\in[d]$ and the perturbations of the queries $\nu_{i,j}$ for each $j\in[d]$, $i\in[r(j)]$. We will fix the random variables $\nu_{i,j}$ for each $j\in[d]$ and $i<r(j)$ by simply conditioning on their randomness; for neighboring graphs, we have a coupling between corresponding variables. For simplicity, we omit this from notation. The remaining random variables are $\hat{T}_{j}$ and $\nu_{r(j),j}$ for each $j\in[d]$; we will be taking probabilities over this randomness, but this will remain implicit in our notation. The main observation needed is that the output of $\mathcal{A}$ is uniquely determined by the iterations $r(j)$ of the first time we have $f_{i,j}(G)+\nu_{i,j}\geq\hat{T}_{j}$ for each $j\in[d]$. Thus, we can define the (deterministic, due to conditioning on the probabilities) quantity
@@
$f_{r(j),j}(G)+\nu_{r(j),j}\geq\hat{T}_{j}\text{ and }\hat{T}_{j}>g_{j}(G)\text{ for each }j\in[d].$ (1)
@@
Since $G,G^{\prime}$ were arbitrary edge-neighboring graphs, this completes the proof. ∎
3.1 Local Privacy Guarantees
In our algorithms, we apply the multidimensional AboveThreshold mechanism in the special case where the queries $\vec{f}_{i}$ have $d=n$ coordinates, and each coordinate of the query corresponds to a node $u$ in the graph $G$ and only depends on the edges $e=(u,v)$ for $v\in V-\{u\}$. In other words, we have an instance of the AboveThreshold on each node $u\in V$, where the data used for the AboveThreshold instance is only the local (edge) data of $u$ [^3] Note that in this setting, an intuitive way to see the privacy guarantee is to view the multidimensional mechanism as a parallel composition of $n$ standard AboveThreshold mechanisms. Since each edge $(u,v)$ is only used in two of the AboveThreshold mechanisms, this implies $2\varepsilon$-differential privacy by parallel composition. However, this is not immediately correct since the queries $\vec{f}_{i}$ are adaptive on the output of all $n$ coordinates of the previous queries, which is not allowed in standard parallel composition. This would require a concurrent version of parallel composition, which does not yet exist.. We will show that in this setting, the multidimensional AboveThreshold mechanism can be implemented locally to satisfy local edge-differential privacy.
For clarity of notation, we index the coordinates of the queries and threshold vector by nodes $v\in V$ instead of indices $j\in[n]$. That is, each query consists of $\vec{f}_{i,v}$ for $v\in V$ and the threshold vector consists of $\vec{T}_{v}$ for $v\in V$. We will now present the changes needed for the local implementation. In Lines 1–3, let each node $u\in V$ compute and store its noisy threshold $\hat{T}_{u}$ using the public threshold $\vec{T}_{u}$. Then for each query $\vec{f}_{i}$ in lines 5–15, each node $u\in V$ can sample its own noise $\nu_{i,u}$ and check the condition $f_{i,u}(G)+\nu_{i,u}\geq\hat{T}_{u}$. The pseudocode is given in Algorithm 2, and we now show that this is locally edge-differentially private.
Input: Private graph $G$, adaptive queries $\{\vec{f}_{i}\}$, threshold vector $\vec{T}$, privacy $\varepsilon$, sensitivity $\Delta$.
1:  for $v\in V$ do
2:     Node $v$ computes $\hat{T}_{v}\leftarrow\vec{T}_{v}+\text{Lap}(2\Delta/\varepsilon)$ and stores it
3:  end for
4:
5:  for each query $\vec{f}_{i}$ do
6:     for $v\in V$ do
7:        Node $v$ samples $\nu_{i,v}\leftarrow\text{Lap}(4\Delta/\varepsilon)$
8:        if $f_{i,v}(G)+\nu_{i,v}\geq\hat{T}_{v}$ then
9:           Node $v$ outputs $a_{i,j}=\top$
10:           Node $v$ outputs STOP, and stops answering queries
11:        else
12:           Node $v$ outputs $a_{i,v}=\bot$
13:        end if
14:     end for
15:  end for
Algorithm 2 Local Multidimensional AboveThreshold
Theorem 3.2.
Algorithm 2 is $\varepsilon$-locally edge differentially private.
Proof.
First, we need to show that the local randomizers are in fact edge-differentially private. This can be seen because each randomizers’ output is a subset of the output of Algorithm 1. In other words, the output is a post-processing of Algorithm 1, so the privacy guarantees follow from Theorem 3.1 and Lemma 2.4. Next, we argue that the transcript is also $\varepsilon$-edge differentially private. Recall that the transcript consists of the set of parties chosen, set of local randomizers assigned, set of privacy parameters assigned, and the set of outputs. In the algorithm, the set of parties chosen, set of local randomizers assigned, and set of privacy parameters assigned are all functions of the outputs from the previous rounds and the public information. Thus, it suffices to prove that the outputs are differentially private by post-processing (Lemma 2.4). But this was already proven in Theorem 3.1, so we are done. ∎
@@ 4 Differentially Private $k$-Core Decomposition
In this section, we give our improved algorithm for differentially private $k$-core decomposition. We first present a variant of the classical (non-private) algorithm for $k$-core decomposition in Section 4.1, and then show how to make it private in Section 4.2. Finally, we show how to implement it in near-linear time in Section 4.3, while incurring only $(1+\eta)$-multiplicative error.
@@ 4.1 A Variation of the Classical Algorithm
Algorithm 3 Threshold-Based $k$-core Decomposition Algorithm
@@ 4.2 Private Implementation of the Algorithm
11:           $\nu_{t,v}\leftarrow\text{Lap}(8/\varepsilon)$
12:           if $d_{V_{t-1}}(v)+\nu_{t,v}\leq k+\tilde{\ell}(v)$ then
@@
Algorithm 4 $\varepsilon$-Differentially Private $k$-Core Decomposition
@@
For each iteration $t$, the $t^{th}$ query consists of the vector of $k-d_{V_{t-1}}(v)$ for each $v\in V$, where $V_{0}=V$ as in the algorithm; note that this matches with the queries on Line 12 of the algorithm. First, observe that the sensitivity of the vector containing $k-d_{V_{t-1}}(v)$ for each $v\in V$ is $\Delta=2$ for each iteration $t$, as needed in the algorithm. Next, observe that we can construct $V_{t}$ from $V_{t-1}$ using only the outputs $a_{t}(v)$ of the queries at the current iteration: given $v\in V_{t-1}$, we include $v$ in $V_{t}$ as well if and only if $a_{t}(v)=\bot$, which is equivalent to the condition in Line 12 of $d_{V_{t-1}}(v)+\nu_{t,v}\leq k+\tilde{\ell}(v)$ [^4] Here, we implicitly assume that the randomness used by the algorithm and the AboveThreshold mechanism are the same. This is justified by a coupling of the random variables in the two algorithms. Specifically, we couple the noise added in Line 4 of the algorithm with the noise added in Line 2 of the AboveThreshold mechanism and couple the noise $\nu_{t,v}$ with $\nu_{i,j}$ in the AboveThreshold mechanism. It is easy to see that for $\Delta=2$, the random variables are exactly the same so the privacy guarantees of multidimensional AboveThreshold translates to privacy guarantees for our algorithm.. Thus, this is a feasible sequence of adaptive queries for the AboveThreshold mechanism.
@@
By the density function of the Laplace distribution, we know that we have $|\nu_{t,u}|\leq\frac{40\log{n}}{\varepsilon}$ and |~$|\tilde{\ell}(u)|\leq\frac{20\log{n}}{\varepsilon}$ each with probability at least $1-\frac{1}{n^{5}}$. Since there are at most $O(n^{3})$ such random variables, taking a union bound over all nodes $u\in V$ and all iterations of the loops, we have the above guarantee with probability at least $1-{O}\left(\frac{1}{n^{2}}\right)$. We condition on the event that the above inequalities hold true for each $t\in[T]$ and each $u\in V$ for the remainder of the proof.
Fix an arbitrary iteration $k$ of the while loop starting on Line 7. Let $H$ be the set of nodes remaining in $V_{t}$ at the end of the while loop in Lines 8—16. We claim that ($i$) all nodes $u\in H$ have core number at least $k-\frac{60\log{n}}{\varepsilon}$ and ($ii$) all nodes $u\not\in H$ have core number at most $k+\frac{60\log{n}}{\varepsilon}$. To see ($i$), consider the subgraph $H$; we claim that each node $u\in H$ has induced degree at least $k-\frac{60\log{n}}{\varepsilon}$ in $H$. Indeed, since each node in $H$ was not removed in the final iteration of the while loop in Lines 8—16, we have that $d_{H_{t}}(u)+\nu_{t,u}\geq k+\tilde{\ell}(u)$ for each $u\in H$. But since we have assumed $|\nu_{t,u}|\leq\frac{40\log{n}}{\varepsilon}$ and |~$|\tilde{\ell}(u)|\leq\frac{20\log{n}}{\varepsilon}$, our desired bounds follow directly by the triangle inequality. To see ($ii$), let’s suppose for contradiction that the $k$-core value of $u$ is $k(u)>\ell+\frac{60\log{n}}{\varepsilon^{\prime}}$. Then there exists a subgraph $u\in K\subseteq V$ where the induced degree of each node $v\in K$ is $d_{K}(v)\geq k(u)$. But for such a subgraph $K$, the condition in Line 12 will always be true (again, because $|\nu_{t,u}|\leq\frac{40\log{n}}{\varepsilon}$ and |~$|\tilde{\ell}(u)|\leq\frac{20\log{n}}{\varepsilon}$) so $K\subseteq H$. But since $u\in K$, this contradicts the fact that $u\not\in H$.
@@ 4.3 Efficient Implementation of the Algorithm
11:           $\nu_{t,v}\leftarrow\text{Lap}(8/\varepsilon)$
12:           if $d_{V_{t-1}}(v)+\nu_{t,v}\leq k+\tilde{\ell}(v)$ then
@@
13:     $t\leftarrow t+1$, $V_{t}\leftarrow V_{t-1}$//this is only notational; we only store the set $V_{t}$ for a single $t$
@@
5 An Improved Analysis of [${}^{+}$22]
We use our MAT technique above to improve the algorithm of [${}^{+}$22] so that we obtain an $\varepsilon$-LEDP algorithm for $k$-core decomposition with additive error $O\left(\frac{\log n}{\varepsilon}\right)$ while maintaining the original $O(\log^{2}n)$ round complexity. We reframe Algorithm 3 of [${}^{+}$22] within our MAT framework in the pseudocode given in Algorithm 7. We highlight the changes we make to the original algorithm in [${}^{+}$22] in blue. Using the improved analysis for the MAT framework, we can obtain our desired error bounds.
Input: Graph $G=(V,E)$, privacy parameter $\varepsilon>0$.
Output: $(2,c^{\prime}\log(n)/\varepsilon)$-approximate core numbers of each node in $G$ for sufficiently large constant $c^{\prime}>0$
1:  Set $\psi=0.1\eta$ and $\lambda=\frac{2(30-\eta)\eta}{(\eta+10)^{2}}$.
2:  for $i=1$ to $n$ do
3:     ~$\tilde{\ell}(i)\leftarrow\text{Lap}(4/\varepsilon)$.
4:  end for
5:  for $r=0$ to $14\log^{2}n-$ 1 do
6:     $L_{r+1}[i]\leftarrow L_{r}[i]$.
7:     if $L_{r}[i]=r$ then
8:        Let $\mathcal{U}_{i}$ be the number of neighbors $j\in\mathbf{a}_{i}$ where $L_{r}[j]=r$.
9:        Sample $X\leftarrow\text{Lap}(8/\varepsilon)$.
10:        Compute $\widehat{\mathcal{U}}_{i}\leftarrow\mathcal{U}_{i}+X$.
11:        if $\widehat{\mathcal{U}}_{i}\leq(1+\psi)^{\lfloor r/(2\log n)\rfloor}+\tilde{\ell}(i)$ then
12:            Stop and set $A_{i}\leftarrow 0$.
13:        else
14:            Continue and set $A_{i}\leftarrow 1$.
15:        end if
16:     end if
17:     $i$ releases $A_{i}$.
18:     $L_{r+1}[i]\leftarrow L_{r}[i]+1$ for every $i$ where $A_{i}=1$.
19:  end for
20:  Curator publishes $L_{r+1}$.
21:  $C\leftarrow\emptyset$.
22:  for $i=1$ to $n$ do
23:     Let $\ell^{\prime}$ be the highest level where $L_{\ell^{\prime}}[i]=\ell^{\prime}$ or 00 if level satisfies this condition.
24:     $\hat{k}(i)\leftarrow(2+\lambda)(1+\psi)^{\max\left(\lfloor\frac{\ell^{\prime}+1}{4\lceil\log_{1+\psi}(n)\rceil}\rfloor-1,0\right)}$.
25:     $C\leftarrow C\cup\{\hat{k}(i)\}$.
26:  end for
27:  return $C$.
Algorithm 7 Improved $\varepsilon$-LEDP $k$-Core Decomposition in $O(\log^{2}n)$ Rounds Based on [${}^{+}$22]
We first show that using our MAT technique given in Section 3 our algorithm is $\varepsilon$-LEDP.
Lemma 5.1.
Algorithm 7 is $\varepsilon$-local edge differentially private.
Proof.
We formulate Algorithm 7 in the context of Algorithm 2 which implies that Algorithm 7 is $\varepsilon$-LEDP. As in the proof of Theorem 4.2, we use an instance of the MAT mechanism with input graph $G$, privacy parameter $\varepsilon$, and threshold vector $\overrightarrow{T}$. We now adaptively define the queries.
For each round $r$, the $r$-th query consists of $(1+\psi)^{\lfloor r/(2\log n)\rfloor}-U_{i}$. The sensitivity of the vector containing such queries for each $v\in V$ is $2$ for each round $r$ (so $\Delta=2$). Let $a_{r}(i)$ be the answers to the adaptive queries from round $r$ for node $i$. Then, we can compute $A_{i}$ (Lines 12 and 14) (via post-processing) for each node $i$ using only $a_{r}(i)$ and $L_{r+1}[i]$ is computed from $A_{i}$ and inductively from $L_{r}[i]$. The approximate core numbers can be obtained from the $L_{r}[i]$ values from each round via post-processing as shown in the original proof in [${}^{+}$22]. Hence, our algorithm can be implemented using MAT and by Theorem 3.2 and Lemma 2.4 is $\varepsilon$-LEDP. ∎
Theorem 5.2.
Algorithm 7 outputs $\left(2+\eta,O\left(\frac{\log n}{\varepsilon}\right)\right)$-approximate core numbers $\hat{k}(v)$ with probability at least $1-O\left(\frac{1}{n^{c}}\right)$ for any constant $c\geq 1$.
We defer the proof of Theorem 5.2 to Appendix A since it is nearly identical to the proof of the original algorithm.
Using the MAT technique in Algorithm 1, we can also remove a factor of $O(\log n)$ in the additive error using the $O(\log n)$ algorithm given in Algorithm 3 of [${}^{+}$22]; however, $\Delta=2\log n$ for that algorithm when reframed in the MAT framework. Thus, we would obtain a $O(\log n)$ round algorithm which is $\varepsilon$-LEDP but gives $O\left(2+\eta,O\left(\frac{\log^{2}n}{\varepsilon}\right)\right)$-approximate core numbers.
@@ 6.1 Differentially Private Densest Subgraph
Algorithm 8 outputs an $\varepsilon$-(local) edge differentially private $\left(2,O\left(\frac{\log(n)}{\varepsilon}\right)\right)$-approximate densest subgraph with probability $1-O\left(\frac{1}{n^{c}}\right)$ for any constant $c\geq 1$.
@@ 6.2 Differentially Private Low Out-Degree Ordering
[${}^{+}$22] defines the low out-degree ordering problem to be finding a differentially private ordering where the out-degree of any vertex is minimized when edges are oriented from vertices earlier in the ordering to later in the ordering. Specifically, there exists an ordering of the vertices of any input graph where the out-degree is at most the degeneracy $d$ of the graph; the degeneracy of the graph is equal to its maximum $k$-core value. The previous best differentially private algorithm for the problem gives out-degree at most $(4+\eta)d+O\left(\frac{\log^{3}n}{\varepsilon}\right)$. In this paper, we give a novel algorithm that obtains out-degree bounded by $d+O\left(\frac{\log n}{\varepsilon}\right)$, with high probability.
Our algorithm (given in Algorithm 9 in Appendix B due to space constraints) is a simple modification of Algorithm 4 where each time we remove a vertex $v$ in Line 13 of Algorithm 4, we add $v$ to the end of our current ordering. The changes to Algorithm 4 are highlighted in blue. Thus, our ordering is precisely the order by which vertices are removed. We now prove that our algorithm for low out-degree ordering is $\varepsilon$-edge DP and gives an ordering for which the out-degree is upper bounded by $d+O\left(\frac{\log n}{\varepsilon}\right)$.
Theorem 6.2.
Our modified Algorithm 4 is $\varepsilon$-(local) edge differentially private and gives a low out-degree ordering where the out-degree is upper bounded by $d+O\left(\frac{\log n}{\varepsilon}\right)$.
Proof.
We first prove that our modified algorithm is $\varepsilon$-(local) edge differentially private. First, the proof of Theorem 4.3 ensures that the ratio of the probabilities that all vertices $v$ are removed at the same steps in edge-neighboring graphs $G$ and $G^{\prime}$ is upper bounded by $e^{\varepsilon}$. Thus, since the ordering of the vertices follows precisely the order by which vertices are removed, then, the ordering that is released is $\varepsilon$-(local) edge DP.
Appendix A Proof of the Approximation Factor in Section 5
Proof of Theorem 5.2.
This proof is nearly identical to the proof of Theorem 4.7 in [${}^{+}$22] except for two details. First, in Algorithm 7, we use Laplace noise instead of symmetric geometric noise and second, we add additional noise to the threshold when determining whether vertices move up levels. Notice that these two changes only affects Invariants 1 and 2. Thus, we only need to show that modified versions of Invariants 1 and 2 still hold for our algorithm and Theorem 4.7 and our approximation proof follow immediately.
Invariant A.1 (Degree Upper Bound).
If node $i\in V_{r}$ (where $V_{r}$ is the set of nodes in level $r$) and level $r<4\log^{2}n-1$, then $i$ has at most $(1+\psi)^{\lfloor r/(2\log n)\rfloor}+\frac{c\log n}{\varepsilon}$ neighbors in levels $\geq r$ whp for large enough constant $c>0$.
Invariant A.2 (Degree Lower Bound).
If node $i\in V_{r}$ (where $V_{r}$ is the set of nodes in level $r$) and level $r>0$, then $i$ has at least $(1+\psi)^{\lfloor(r-1)/(2\log n)\rfloor}-\frac{c\log n}{\varepsilon}$ neighbors in levels $\geq r-1$ whp for large enough constant $c>0$.
First, we know that by the density function of the Laplace distribution, any noise drawn from the Laplace distributions used in our algorithm is upper bounded by $\frac{c\log{n}}{\varepsilon}$ for large enough constant $c>0$ whp. We first show Invariant A.1. A node moves up a level when $\widehat{\mathcal{U}}_{i}$ exceeds the threshold given by $(1+\psi)^{\lfloor r/(2\log n)\rfloor}+\tilde{\ell}(i)$. By our algorithm, $\widehat{\mathcal{U}}_{i}=\mathcal{U}_{i}+X$ where $X$ and ~$\tilde{\ell}(i)$ are noises drawn from the Laplace distribution. Thus, if a node doesn’t move up, then $\mathcal{U}_{i}\leq(1+\psi)^{\lfloor r/(2\log n)\rfloor}+\tilde{\ell}(i)-X$. The expression on the right is at most $(1+\psi)^{\lfloor r/(2\log n)\rfloor}+2\cdot\frac{c\log{n}}{\varepsilon}$ whp for large enough constant $c>0$.
Now, we prove Invariant A.2. If a node is on level $>0$, then it must have moved up at least one level. If it moved up at least one level (from level $r-1$), then its $\widehat{\mathcal{U}}_{i}$ at level $r-1$ must be at least $(1+\psi)^{\lfloor(r-1)/(2\log n)\rfloor}+\tilde{\ell}(i)$. Then, it holds that $\mathcal{U}_{i}\geq(1+\psi)^{\lfloor(r-1)/(2\log n)\rfloor}+\tilde{\ell}(i)-X$ at level $r-1$. The right hand side is at least $(1+\psi)^{\lfloor(r-1)/(2\log n)\rfloor}-2\cdot\frac{c\log n}{\varepsilon}$ for large enough constant $c>0$ and Invariant A.2 also holds.
We now can use the proof of Theorem 4.7 exactly as written except with the new bounds given in Invariant A.1 and Invariant A.2 to obtain our final approximation guarantee. ∎
Appendix B Algorithm for Low Out-Degree Ordering